# A Unified Framework for Model Editing

### v2

@@ Abstract
We introduce a unifying framework that brings two leading "locate-and-edit" model editing techniques – ROME and MEMIT – under a single conceptual umbrella, optimizing for the same goal, which we call the preservation-memorization objective. ROME uses an equality constraint to perform one edit at a time, whereas MEMIT employs a more flexible least-square constraint that allows for batched edits. Following the preservation-memorization objective, we present Equality-constrained Mass Model Editing algorithm for Transformers or EMMET, a new batched memory-editing algorithm that uses a closed-form solution for the equality-constrained version of the preservation-memorization objective. EMMET is a batched-version of ROME and is able to perform batched-edits up to a batch-size of 10,000 with very similar performance to MEMIT across multiple dimensions. With EMMET, we unify and achieve symmetry within the "locate-and-edit" algorithms, allowing batched-editing using both objectives.
@@ 1 Introduction
As new facts emerge constantly, it’s crucial to keep models up-to-date with the latest knowledge. Model editing Yao et al. (2023) gives us the ability to edit facts stored inside a model as well as update incorrectly stored facts. In this paper, we focus on two popular parameter modifying model editing methods which infuse knowledge within models without needing an additional hypernetwork Chauhan et al. (2023). These methods are ROME (Rank-One Model Editing) (Meng et al., 2022a) and MEMIT (Mass Editing Memory in Transformer) (Meng et al., 2022b). These methods directly update specific "knowledge-containing" parts of the model without requiring the need to train additional models and can be applied to any transformer based large language model (LLMs). MEMIT also uniquely allows for batched edits (appendix A.1).
@@
In this paper, we present a unifying conceptual framework for ROME and MEMIT and show that both methods optimize the same objective. We call this the preservation-memorization objective of model editing, where new knowledge is injected or memorized such that representations of certain vectors are preserved through the editing process. We show that ROME optimizes an equality-constrained version of the objective whereas MEMIT optimizes a more relaxed least-squares version of the objective, which allows for a simple closed-form solution for making batched edits. We then highlight that MEMIT consists of two separate steps - an optimization objective and an algorithm that distributes the edits into multiple layers. The power of MEMIT in many cases comes from these edit-distribution algorithms.
Finally, we present a closed-form solution for making batched edits with the equality-constraint under the preservation-memorization objective in the form of EMMET - an Equality-constrained Mass Model Editing algorithm for Transformers. With EMMET, batched edits can be performed for batch sizes upto 10,000 with its performance matching MEMIT across multiple dimensions. We evaluate EMMET on three models - GPT2-XL Radford et al. (2019), GPT-J Wang and Komatsuzaki (2021) and Llama-2-7b Touvron et al. (2023) on standard model editing datasets - CounterFact Meng et al. (2022a, b) and zsRE Levy et al. (2017). The code for EMMET can be found here [^1] https://github.com/myanonymousrepo/unified_model_editing.
@@
  • We unify two popular model editing techniques (ROME and MEMIT) under the same conceptual framework called the preservation-memorization.
  • We disentangle the MEMIT objective from the MEMIT algorithm which distributes edits within multiple layers. We hope this sparks further research in edit-distribution algorithms.
  • We present a closed-form solution to equality-constrained memorization in the form of EMMET, a batched version of ROME. EMMET is a new batched-editing algorithm that achieves symmetry between the two objectives of "locate-and-edit" class of algorithms and shows that batched edits can be made using both objectives.
@@ 3 Preservation-Memorization : A Unifying Framework for ROME and MEMIT
Let $W$ represent the weights of the feed-forward layer we want to edit [^2] These layers are found by causal tracing methods (Meng et al., 2022a, b), and let $k$ be a key-vector representative of a fact that we are either editing or preserving, and is the input vector to $W$. The layers being edited are shown in an expanded diagram of a transformer layer Vaswani et al. (2017) in Figure 2. The output of a single transformer layer for a general decoder-only LLM can be written as:
$a^{l}$ $={LN}(\texttt{Att}(h^{l-1})+h^{l-1})$ (1)
$m^{l}$ $=W^{l}_{proj}\left({NL}(W^{l}_{fc}a^{l}+b^{l}_{fc})\right)+b^{l}_{proj}$ (2)
$h^{l}$ $={LN}(m^{l}+a^{l})$ (3)
Here, Att is the multi-head attention function, NL refers to the non-linearity used in the MLP module of the model and LN refers to layer normalization. The keys are outputs of the first linear layer, or $k^{l}=\texttt{NL}(a^{l}W^{l}_{fc}+b^{l}_{fc})$, whereas $v^{l}=m^{l}$. A detailed explanation on creation of key-vectors and value-vectors is given in Appendix A.3.
In the model editing process, the weights of an intermediate layer of the model are changed from $W_{0}$ to $\hat{W}$, where $k_{0}$ is used to indicate a key-vector representing facts we want to preserve from the original model, and $k_{e}$ being key-vectors representing facts we want to insert into the model. Let $v_{e}$ be the desired output at the layer being edited corresponding to input $k_{e}$ such that the correct fact is recalled by the model when finally generating text.
@@
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|}_{\text{preservation}}\hskip 4.0pt\text{ s.t. }\underbrace{\hat{W}k_{e}=v_{e}}_{\text{memorization}}$ (4)
@@
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\lambda||\hat{W}K_{0}-W_{0}K_{0}||}_{\text{preservation}}+\underbrace{||\hat{W}K_{E}-V_{E}||}_{\text{memorization}}$ (7)
where $K_{E}=[k^{e}_{1}\hskip 4.0pt|k^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{e}_{E}]$ is a matrix containing a row of vectors representing the edits we are making and $V_{E}=[v^{e}_{1}\hskip 4.0pt|v^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptv^{e}_{E}]$ represents their target representations.
@@ 4 Edit-Distribution Algorithms
Prior works on model editing do not differentiate between the MEMIT-objective and the edit-distribution algorithm, and as a consequence we never see edits using ROME being distributed to multiple layers or MEMIT being used on only a single layer.
@@ 4.1 Impact of edit-distribution Algorithms
The key advantage of the edit-distribution algorithm is apparent when making batched edits. In this section, we perform two experiments to analyze this. First, we compare single edits in ROME and MEMIT with and without edit distribution on 1k randomly selected facts from the CounterFact dataset. Following that, we compare batched editing in MEMIT with and without edit distribution. Both experiments are performed on three different models - GPT2-XL (1.5B), GPT-J (6B) and Llama2-7B.
The results are shown in Table 1 for edits without distribution and Table 3 for edits with distribution. We use the more stable version of ROME called r-ROME as presented in (Gupta and Anumanchipalli, 2024) that does not lead to model collapse and improves generalization. We see that solutions to both ROME and MEMIT objectives perform equally well at making singular edits across different metrics, without needing to distribute the edits to multiple layers. To highlight the usefulness of edit-distribution algorithms, we compare MEMIT when making batched edits. The results are shown in Figure 3. When only editing a single layer, we see that MEMIT is able to successfully make batched edits up to a batch size of 1024 for GPT2-XL, 256 for Llama-2-7b and a batch-size as large as 4096 for GPT-J [^3] In our experiments we find GPT-J to be an easier model to edit compared to other models. This is both intriguing but also not the best model to evaluate model editing success.. After this point, the performance of model editing increases when making edits on multiple layers, except for Llama-2-7b. All hyperparameters for all models were chosen as is from prior work (Meng et al., 2022a, b; Yao et al., 2023; Zhang et al., 2024) (appendix A.2).
With these experiments, we want to highlight two key points - firstly, when comparing the effectiveness of two optimization objectives, the evaluation should not be conflated with the edit distribution algorithms. Secondly, the MEMIT edit-distribution algorithm is not perfect and currently is the only way to distribute edits into multiple layers, where the residual in the update is distributed with specific ratios through different layers. We hope these experiments will bring more focus to edit distribution algorithms and boost further research in these methods.
@@ 5 Introducing EMMET
As can be seen in the above equation, the preservation of representations happens in the first term whereas memorization of all the new facts are forced using an equality constrain in the second term. The above equation is solved using lagrange-multipliers. The Lagrangian for the above equation for multiple equality constraints requires a summation of lagrange multipliers and equals:
$\hat{W}K_{0}K_{0}^{T}=W_{0}K_{0}K_{0}^{T}+\sum_{i=1}^{E}\lambda_{i}k^{e^{T}}_{i}\$ (13)
which is same as:
@@
$\hskip 8.0pt\hat{W}=W_{0}+\Lambda K_{E}^{T}C^{-1}_{0}\$ (15)
@@
We write the update equation of EMMET in a familiar form, where the residual $R=V_{E}-W_{0}K_{E}$ is modified by some matrix operations to update the models with new edits. Additionally, when we put $E=1$, the $K_{E}$ matrix reduces to a single vector $k_{e}$ and equation 17 reduces to the ROME update equation (equation 5). With EMMET, we complete the unification of ROME and MEMIT under the preservation-memorization objective and achieve a symmetry with the usage of these algorithms. EMMET allows for making batched-edits as well as singular when using equality constraints for memorization, much similar to MEMIT with least-square based memorization.
@@ 5.1 Stabilizing EMMET
where $k^{0}_{i}$ represent a key-vector we want to preserve. For an LLM of dimension $d$, the dimensionality of a key-vector is usually $4d$ (Figure 2), which is the dimensionality of the square matrix $C_{0}$. If $C_{0}$ is a $4d$-dimensional square matrix which is a summation of rank-1 matrices, it is invertible as long as there are atleast $4d$-independent vectors in the summation, or $4d$-independent vectors in $K_{0}$. For example, for Llama-2-7b with hidden dimension of 4096, the dimensionality of key vectors are 16384. So as long as representations of atleast 16384 independent key-vectors are being preserved while editing, $C_{0}$ will be an invertible matrix. In practice, we preserve representations of a much larger number of vectors, and hence this condition is almost always satisfied.
The matrix $D=K_{E}^{T}C^{-1}_{0}K_{E}$ is a square matrix of dimensionality equal to the number of edits. If given that $C_{0}$ is invertible, $D$ is invertible as long as $K_{E}$ is full-rank, which means all key-vectors corresponding to facts being memorized are independent of each other. While this is not guaranteed, it can be verified before editing and facts corresponding to non-independent keys can be removed from a batch. In practice, we do not find invertibility of D being an issue. However, we find that $D$ is often ill-conditioned, which means that the ratio of the largest and smallest eigenvalues of $D$ explodes. This doesn’t necessarily mean that the matrix is singular (non-invertible), but it does mean that numerical computations involving the matrix inverse are unstable and can lead to large numerical errors. To counter this, we set $D=D+\alpha I$, where $\alpha$ is set to 0.1 after an ablation over multiple batch sizes. This allows for stable batched edits using EMMET.
@@ 5.2 Batch Editing with EMMET
We begin by experimenting with EMMET for model editing with varied batch sizes on GPT2-XL, GPT-J and Llama-2-7b on the CounterFact and zsRE Levy et al. (2017) datasets. The exact implementation details can be found in section A.2. We compare the performance of EMMET and MEMIT on batch sizes up to 10,000 while editing both single (to directly compare the optimization objectives) and multiple layers. The single layer editing comparison between EMMET and MEMIT can be found in Figure 4. We see that both methods have almost identical performance in practice across different metrics. MEMIT performs slightly better than EMMET for Llama-2-7b, as indicated by ES, PS and S metrics. We then apply the MEMIT edit-distribution on EMMET and compare it with MEMIT. The results are shown in Figure 5. We see that in this case, EMMET performs slightly better than MEMIT for Llama-2-7b. The results on the zsRE dataset tell a similar story and can be seen in Figure 7 and 8. These results present EMMET as a viable new batched-editing algorithm.
@@
Previous work (Gu et al., 2024; Gupta et al., 2024) has shown that model editing is often accompanied by model degradation. Gupta et al. (2024) show this by evaluating the edited model on downstream tasks from the popular GLUE benchmark Wang et al. (2018). We adopt their evaluation setting and evaluate both EMMET and MEMIT on four downstream tasks - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019) for doing downstream evaluation. The results are shown in Figure 6 for a batch size of 256. The results for other batch sizes can be found in Appendix A.2. We find that both EMMET and MEMIT also degrade the model similarly.
Although EMMET is unable outperform MEMIT, it is an important piece in unifying model editing under the preservation-memorization framework. Both algorithms are able to make successful batched edits upto a batch-size of 10,000 and lead to similar model degradation. EMMET imposes a "theoretically" stronger memorization constraint, yet we do not see an improvement in editing efficacy. This indicates that we may be reaching the limit of model editing capabilities under the preservation-memorization objective.

### v4

@@ Abstract
ROME and MEMIT are largely believed to be two different model editing algorithms, with the major difference between them being the ability to perform batched edits. In this paper, we unify these two algorithms under a single conceptual umbrella, optimizing for the same goal, which we call the preservation-memorization objective. ROME uses an equality constraint to optimize this objective to perform one edit at a time, whereas MEMIT employs a more flexible least-square constraint that allows for batched edits. We generalize ROME and enable batched editing with equality constraint in the form of EMMET - an Equality-constrained Mass Model Editing algorithm for Transformers, a new batched memory-editing algorithm. EMMET can perform batched-edits up to a batch-size of 10,000, with very similar performance to MEMIT across multiple dimensions. With the introduction of EMMET, we truly unify ROME and MEMIT and show that both algorithms are equivalent in terms of their optimization objective, their abilities (singular and batched editing), their model editing performance and their limitations.
@@ 1 Introduction
As new facts emerge constantly, it is crucial to keep models up-to-date with the latest knowledge. Model editing Yao et al. (2023) gives us the ability to edit facts stored inside a model as well as update incorrectly stored facts. In this paper, we focus on two of the most popular and best performing model editing methods - ROME (Rank-One Model Editing) (Meng et al., 2022a) and MEMIT (Mass Editing Memory in Transformer) (Meng et al., 2022b). ROME and MEMIT directly update specific "knowledge-containing" parts of the model without requiring the need to train additional models (De Cao et al., 2021; Mitchell et al., 2021; Tan et al., 2023) and can be applied to any transformer based large language model (LLMs). This makes these algorithms really attractive for practical use cases. MEMIT also uniquely allows for batched edits (appendix 5.3).
@@
ROME and MEMIT are largely considered different from each other, with one of their major differences being that ROME allows for editing only one fact at a time. In this paper, we present a unifying conceptual framework for ROME and MEMIT and show that both methods optimize the same objective function. We call this the preservation-memorization objective of model editing, where new knowledge is injected or memorized such that representations of certain vectors are preserved through the editing process. We show that ROME optimizes an equality-constrained version of the objective whereas MEMIT optimizes a more relaxed least-squares version of the objective, which allows for a simple closed-form solution for making batched edits. We then highlight that MEMIT consists of two separate steps - an optimization objective and an algorithm that distributes the edits into multiple layers. The power of MEMIT in many cases comes from these edit-distribution algorithms.
Finally, we present a closed-form solution for making batched edits with equality-constraint under the preservation-memorization objective in the form of EMMET - an Equality-constrained Mass Model Editing algorithm for Transformers. With EMMET, batched edits can be performed for batch sizes up to 10,000 with performance much similar to MEMIT. We evaluate EMMET on three models - GPT2-XL, GPT-J and Llama-2-7b on standard model editing datasets - CounterFact and zsRE. Enabling batched editing with equality-constraint in the form of EMMET allows us to truly unify the two algorithms and shows that both ROME and MEMIT are essentially equivalent in terms of their optimization objective, their abilities (performing singular and batched editing), their model editing performance and their limitations. EMMET serves as a cornerstone in completing this larger picture. The code for EMMET can be found here [^1] https://github.com/scalable-model-editing/unified-model-editing.
@@
  • We unify two popular model editing techniques (ROME and MEMIT) under the preservation-memorization objective and show that these algorithms are equivalent in terms of their optimization objective and in practice.
  • We disentangle the MEMIT objective from the MEMIT algorithm which distributes edits within multiple layers. This allows for a fair comparison of MEMIT and ROME.
  • We present a closed-form solution to equality-constrained memorization in the form of EMMET, a batched version of ROME. EMMET is a new batched-editing algorithm that achieves symmetry in usage and performance between the two algorithms and shows that batched edits can be made using both objectives.
@@ 3 Preservation-Memorization : A Unifying Framework for ROME and MEMIT
Let $W$ represent the weights of the feed-forward layer we want to edit [^2] These layers are found by causal tracing methods (Meng et al., 2022a, b), and let $k$ be a key-vector representative of a fact that we are either editing or preserving, and is the input vector to $W$. The layers being edited are shown in an expanded diagram of a transformer layer Vaswani et al. (2017) in Figure 2. In the model editing process, the weights of an intermediate layer of the model are changed from $W_{0}$ to $\hat{W}$ ($W_{0}$ represents the original weights of the $W_{proj}$ matrix), where $k_{0}$ is used to indicate a key-vector representing facts we want to preserve from the original model, and $k_{e}$ being key-vectors representing facts we want to insert into the model. Let $v_{e}$ be the desired output at the layer being edited corresponding to input $k_{e}$ such that the correct fact is recalled by the model when finally generating text. A detailed explanation on creation of key-vectors and value-vectors is given in Appendix A.2 and is also briefly depicted in Figure 1.
@@
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|^{2}_{F}}_{\text{preservation}}\hskip 4.0pt\text{ s.t. }\underbrace{\hat{W}k_{e}=v_{e}}_{\text{memorization}}$ (1)
@@
$\underset{\hat{W}}{\operatorname{argmin}}\hskip 4.0pt\underbrace{\lambda\left\|\hat{W}K_{0}-W_{0}K_{0}\right\|^{2}_{F}}_{\text{preservation}}+\underbrace{\left\|\hat{W}K_{E}-V_{E}\right\|^{2}_{F}}_{\text{memorization}}$ (4)
where $K_{E}=[k^{e}_{1}\hskip 4.0pt|k^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptk^{e}_{E}]$ is a matrix containing a row of vectors representing the edits we are making in a batch and $V_{E}=[v^{e}_{1}\hskip 4.0pt|v^{e}_{2}\hskip 4.0pt|\dots|\hskip 4.0ptv^{e}_{E}]$ represents their target representations.
@@ 4 Edit-Distribution Algorithms
Prior works on model editing do not differentiate between the MEMIT-objective and the edit-distribution algorithm, and as a consequence we never see edits using ROME being distributed to multiple layers or MEMIT being used on only a single layer. The additional wrapping of edit-distribution also makes MEMIT seem distant from ROME. In the next section, we remove the wrapping of edit-distribution from MEMIT and allow for a fair comparison between the two algorithms.
@@ 4.1 Impact of edit-distribution Algorithms
The key advantage of the edit-distribution algorithm is apparent when making batched edits. In this section, we perform two experiments to analyze this. First, we compare single edits in ROME and MEMIT with and without edit distribution on 1k randomly selected facts from the CounterFact datase Meng et al. (2022a). Following that, we compare batched editing in MEMIT with and without edit distribution. Both experiments are performed on three different models - GPT2-XL (1.5B) Radford et al. (2019), GPT-J (6B) Wang and Komatsuzaki (2021) and Llama2-7B Touvron et al. (2023).
The results are shown in Table 1 for edits without distribution and Table 3 (appendix) for edits with distribution. We use the more stable version of ROME called r-ROME as presented in (Gupta and Anumanchipalli, 2024) that does not lead to model collapse and improves generalization. We see that solutions to both ROME and MEMIT objectives perform equally well at making singular edits across different metrics, without needing to distribute the edits to multiple layers. To highlight the usefulness of edit-distribution algorithms, we make batched edits with MEMIT comparing performance with and without edit distribution. The results are shown in Figure 3. When only editing a single layer, we see that MEMIT is able to successfully make batched edits up to a batch size of 1024 for GPT2-XL, 256 for Llama-2-7b and a batch-size as large as 4096 for GPT-J [^3] In our experiments we find GPT-J to be an easier model to edit compared to other models. This is both intriguing but also not the best model to evaluate model editing success.. After this point, the performance of model editing increases when making edits on multiple layers, except for Llama-2-7b. All hyperparameters for all models were chosen as is from prior work (Meng et al., 2022a, b; Yao et al., 2023; Zhang et al., 2024) (appendix A.1).
With these experiments, we want to highlight two key points - firstly, when comparing the effectiveness of two optimization objectives, the evaluation should not be conflated with the edit distribution algorithms. After removing the wrapping of edit-distribution from MEMIT, we see that the performance numbers for ROME and MEMIT have an uncanny similarity. Secondly, the MEMIT edit-distribution algorithm is not perfect and currently is the only way to distribute edits into multiple layers, where the residual in the update is distributed with specific ratios through different layers. We hope these experiments will bring more focus to edit distribution algorithms and boost further research in these methods.
@@ 5 Introducing EMMET
As can be seen in the above equation, the preservation of representations happens in the first term whereas memorization of all the new facts are forced using an equality constraint in the second term. The above equation is solved using lagrange-multipliers. The derivation of the above equation for the generalized case of batched editing can be found in Appendix A.3.
@@
The closed form solution for batched editing with equality-constraint or EMMET is shown below:
@@
Here, $C_{0}=K_{0}K_{0}^{T}$ has the usual meaning as in the derivation of ROME and MEMIT, where $K_{0}$ contains the list of representations we want preserved during editing. We write the update equation for EMMET in a familiar form, where the residual $R=V_{E}-W_{0}K_{E}$ is modified by some matrix operations to update the models with new edits. Additionally, when we put $E=1$, the $K_{E}$ matrix reduces to a single vector $k_{e}$ and equation 9 reduces to the ROME update equation (equation 2). With EMMET, we complete the unification of ROME and MEMIT under the preservation-memorization objective and achieve a symmetry with the usage of these algorithms. EMMET allows for making batched-edits as well as singular when using equality constraints for memorization, much similar to MEMIT with least-square based memorization.
@@ 5.1 Stabilizing EMMET
where $k^{0}_{i}$ represents a key-vector we want to preserve. For an LLM of dimension $d$, the dimensionality of a key-vector is usually $4d$ (Figure 2), which is the dimensionality of the square matrix $C_{0}$. If $C_{0}$ is a $4d$-dimensional square matrix which is a summation of rank-1 matrices, it is invertible as long as there are atleast $4d$-independent vectors in the summation, or $4d$-independent vectors in $K_{0}$. For example, for GPT2-XL with hidden dimension of 1600, the dimensionality of key vectors are 6400. So as long as representations of atleast 6400 independent key-vectors are being preserved while editing, $C_{0}$ will be an invertible matrix. In practice, we preserve representations of a much larger number of vectors, and hence this condition is always satisfied.
The matrix $D=K_{E}^{T}C^{-1}_{0}K_{E}$ is a square matrix of dimensionality equal to the number of edits. If given that $C_{0}$ is invertible, $D$ is invertible as long as $K_{E}$ is full-rank, which means all key-vectors corresponding to facts being memorized are independent of each other. While this is not guaranteed, it can be verified before editing and facts corresponding to non-independent keys can be removed from a batch. In practice, we do not find invertibility of D being an issue. However, we find that $D$ is often ill-conditioned, which means that the ratio of the largest and smallest eigenvalues of $D$ explodes. This doesn’t necessarily mean that the matrix is singular (non-invertible), but it does mean that numerical computations involving the matrix inverse are unstable and can lead to large numerical errors. To counter this, we set $D=D+\alpha I$, where $\alpha$ is set to 0.1 after an ablation over multiple batch sizes. This allows for stable batched edits using EMMET and also ensures that the $D$ matrix is always invertible.
@@ 5.2 Batch Editing with EMMET
We begin by experimenting with EMMET for model editing with varied batch sizes on GPT2-XL, GPT-J and Llama-2-7b on the CounterFact and zsRE Levy et al. (2017) datasets. The exact implementation details can be found in section A.1. We compare the performance of EMMET and MEMIT on batch sizes up to 10,000 while editing both single (to directly compare the optimization objectives) and multiple layers. The single layer editing comparison between EMMET and MEMIT can be found in Figure 4. We see that both methods have almost identical performance in practice across different metrics. MEMIT performs slightly better than EMMET for Llama-2-7b, as indicated by ES, PS and S metrics. We then apply the MEMIT edit-distribution on EMMET and compare it with MEMIT. The results are shown in Figure 5. We see that in this case, EMMET performs slightly better than MEMIT for Llama-2-7b. The results on the zsRE dataset tell a similar story and can be seen in Figure 7 and 8. The experiments for different hyperparameter values are shown in Appendix A.4. These results present EMMET as a viable new batched-editing algorithm.
@@
Previous work (Gu et al., 2024; Gupta et al., 2024) has shown that model editing is often accompanied by model degradation. This was shown by evaluating the edited model on downstream tasks from the popular GLUE benchmark Wang et al. (2018). Once we identified that memorization in MEMIT is happening using an approximate least-square constraint rather than an equality constraint, we hypothesised that a possible reason for model degradation could be the use of the least-square constraint. Thus, using an equality constraint, which by definition requires the edit to be exact, may not degrade other knowledge or skills of the model. This was also the motivation behind generalizing ROME to batched edits in the form of EMMET. To test this hypothesis, we adopt the evaluation setting of Gupta et al. (2024) and evaluate both EMMET and MEMIT on four downstream tasks - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019) for doing downstream evaluation. The results are shown in Figure 6 for a batch size of 256. The results for other batch sizes can be found in Appendix A.1. We find that both EMMET and MEMIT also degrade the model similarly.
The fact that both EMMET and MEMIT perform editing and degrade the model with an uncanny similarity shows that a "stronger" equality constraint does not enable more accurate model editing. We believe reason behind this is the construction of the key-vector, which is created by taking the average of representations of multiple phrasings of a fact (appendix A.2). This is done to make edits that generalize beyond a single phrasing of a fact. As the key-vector is an averaged representation over randomly selected phrasings, it is an approximation of the ideal vector representation of a fact. We believe that such an approximate representation does not require the additional accuracy of memorization enforced due to the equality constraint. Our findings also indicate that we may be reaching the limit of model editing capabilities under the preservation-memorization objective.
5.3 Related Work